OpenStudy (anonymous):

Divide eight and two-thirds divided by two and one-third. seven twenty-sixths four and two-thirds three and five-sevenths four and one-half

OpenStudy (anonymous):

"frak" Medals Given

OpenStudy (shamil98):

\[\huge 8\frac{ 2 }{ 3 } \div 2\frac{ 1 }{ 3 } --> \frac{ 26 }{ 3 } \div \frac{ 7 }{ 3 }\] when dividing fractions you multiply by the reciprocal \[\huge \frac{ 26 }{ 3 } \times \frac{ 3 }{ 7 } = \frac{ 26 \times 3 }{ 3 \times 7 }\] multiply and put in mixed fraction form.

OpenStudy (anonymous):

Thank you Sham I can get the rest from here